OpenStudy (anonymous):

solve for x 2/5+3/5x=x+5/10

OpenStudy (diyadiya):

\[\frac{2}{5}+\frac{3}{5}x= x+ \frac{5}{10}\] first subtract 2/5 from both sides

OpenStudy (diyadiya):

what do you get?

OpenStudy (anonymous):

3/5x=x+5/10-2/5

OpenStudy (diyadiya):

Yes btw what's 5/10 - 2/5 ?